A square lawn is surrounded by a path 2.5 m wide around it . If the area of the path is 200 sq m , find the area of the lawn
Answer by mananth(16949)

let the side be x
Area of square = side ^2
area = x^2
width = 2.5
length increases by 5 m
Area including width = (x+5)(x+5)= (x+5)^2
(x+5)^2-x^2=2000
x^2+10x+25-x^2=200
10x=200-25
10x=175
/10
x= 17.5 m